A circuit is constructed with a 1000 ​Ω​ resistor, a 1.00 x 10​-9​ F capacitor and a 2.00 H inductor. This circuit is connected to a 100 V​max​ generator. The frequency of the generator can be varied. What frequency will produce the lowest impedance? What is the rms current in the circuit at this frequency?
Answer #1

Lowest impedence of the circuit will be when inductive impedence and capacitive impedence are equal

Z = R

And,

1) Xc = Xl

F = 1/(2*3.14*sqrt(L*C))

F = 3560.61 Hz

2) Irms = Vmax/1.414*R

Irms = 100/1.414*1000

Irms = 0.070721 A